Exercise 5. [A-M Ch 3 Ex 7] Let R#0 be a ring. A multiplicatively closed subset S of R is said to be saturated if XY ES #xe S1. Let I be the collection of all multiplicatively closed subsets of R such that 0 € S. Show that I has maximal elements, andCan you please provide clear and step by step solution for both 3 and 4. Thanks :)
Answer #1

I @ S = 1+ I,. I ar A prime iseal o murs s if we have at I and xep sit. x=lta on 1-a-x -> <ty = I + P in RF I to Thin the uni